Hi,
I'm having a hard time figuring out the following word problem, and I was hoping you could take the time to explain/ show me how to work it. Thanks in advance! :)
Part of the proceeds from a garage sale was $280 worth of $5 and $10 bills. If there were 20 more $5 bills than $10 bills, find the number of each denomination.
What I got out of it is the following:
Let F= $5 and T= $10
Then F= T+20
10T + 5(T+20)= 280
I'm fustrated and I can't figure out what I'm doing wrong. Thank You!!!

Answer by mananth(16949)

$10.00---------x
$5.00----------5x+20
10x+5x+100=280
15x+100 =280
15x=280-100
5x=180
/15
x=12 $10.00 bills =120
12+20=32----- 5$ 5 bills
32*5= $160